### v1

# AltLoRA: Towards Better Gradient Approximation in Low-Rank Adaptation with Alternating Projections

Abstract
Low-Rank Adaptation (LoRA) has emerged as an effective technique for reducing memory overhead in fine-tuning large language models. However, it often suffers from sub-optimal performance compared with full fine-tuning since the update is constrained in the low-rank space. Recent variants such as LoRA-Pro attempt to mitigate this by adjusting the gradients of the low-rank matrices to approximate the full gradient. However, LoRA-Pro’s solution is not unique, and different solutions can lead to significantly varying performance in ablation studies. Besides, to incorporate momentum or adaptive optimization design, approaches like LoRA-Pro must first compute the equivalent gradient, causing a higher memory cost close to full fine-tuning. A key challenge remains in integrating momentum properly into the low-rank space with lower memory cost. In this work, we propose AltLoRA, an alternating projection method that avoids the difficulties in gradient approximation brought by the joint update design, meanwhile integrating momentum without higher memory complexity. Our theoretical analysis provides convergence guarantees and further shows that AltLoRA enables stable feature learning and robustness to transformation invariance. Extensive experiments across multiple tasks demonstrate that AltLoRA outperforms LoRA and its variants, narrowing the gap toward full fine-tuning while preserving superior memory efficiency.
1 Introduction
Low-Rank Adaptation (LoRA [23]) has emerged as a leading approach for parameter-efficient fine-tuning (PEFT)([22, 36, 33]) of large language models ([4, 48, 57, 38]). Building on prior work investigating the intrinsic dimensionality of neural networks ([2, 34]), LoRA assumes that fine-tuning updates can be effectively captured in a low-rank subspace. Specifically, for a pre-trained model with weight matrix $W_{0}\in\mathbb{R}^{k\times d}$, LoRA reparameterizes the weight update $\Delta W$ via a low-rank decomposition as $W_{0}+\Delta W=W_{0}+sBA,$ where $B\in\mathbb{R}^{k\times r}$, $A\in\mathbb{R}^{r\times d}$ and $s=\frac{\alpha}{r}$ is a scaling factor. Here, $r\ll\min(k,d)$ is the rank of the update. Thanks to its substantial memory and computational savings [23], LoRA has enabled scalable adaptation across diverse applications, including reinforcement learning from human feedback (RLHF) [53, 21], diffusion models [41, 73], and mixture-of-experts (MoE) architectures [63, 35].
Despite its parameter efficiency, LoRA often underperforms full fine-tuning ([12, 23, 39, 67]). This gap has fueled growing interest in optimizing LoRA via hyperparameter tuning under stable feature learning [19, 18] and optimizers that preserve transformation invariance [75]. Formally, if we denote the loss function as $L$, full fine-tuning will utilize the full gradient $\nabla_{W}L\in\mathbb{R}^{k\times d}$ for backpropagation. In contrast, the gradients in LoRA for $B$ and $A$ are given by $(\nabla_{W}L)A^{\top}$ and $B^{\top}(\nabla_{W}L)$, respectively (see Section 3). This reparameterization significantly alters the gradient flow during training [84] by restricting it to the low-rank space.
A promising direction to fill the gap between the gradient dynamics is to ensure that the equivalent gradient established by LoRA approximates the full gradient ([62, 61, 47]). However, two key challenges in the gradient approximation for low-rank adaptation remain unaddressed. First, LoRA-Pro  [62] depends on an auxiliary variable that impacts the performance significantly. Depending on the choice of this variable, the evaluation score varies from $31.74$ to $57.57$ on the GSM8K datasets (see Appendix D.1 in [62]). Obtaining a unique solution requires solving a Sylvester equation, which introduces additional computational cost and relies on a non-standard assumption. Second, as LoRA-Pro accelerates the equivalent gradient with full-parameter learning, it requires a memory cost like full fine-tuning with space complexity $\mathcal{O}(kd)$ as shown in Table 1. In contrast, LoRA maintains a more efficient space complexity of $\mathcal{O}(kr+rd)$. Under such memory constraints, how to incorporate momentum properly within the low-rank structure is largely unexplored.
In this paper, to close the performance gap between LoRA and full fine-tuning, we address the two key challenges outlined above and propose a novel PEFT method, AltLoRA, based on Alternating updates to the Low-Rank Adaptation. AltLoRA properly approximates the full gradient by alternately projecting it onto low-rank subspaces and $B$. Building on this projection-based gradient approximation, we further introduce a new mechanism to optimize momentum effectively within the low-rank space, while strictly adhering to the memory constraints of LoRA [23]. Without allowing full-parameter learning, AltLoRA is the first work in the literature to properly optimize both gradient and momentum over the low-rank subspaces, while achieving stable feature learning and transformation invariance, as summarized in Table 1.
Methods Gradient Approximation Stable Feature Learning Transformation Invariance Time Complexity Space Complexity
LoRA [23] $\mathcal{O}(kr^{2}+dr^{2})$ $\mathcal{O}(kr+dr)$
LoRA+ [19] $\mathcal{O}(kr^{2}+dr^{2})$ $\mathcal{O}(kr+dr)$
ScaledAdam [77] $\mathcal{O}(kr^{2}+dr^{2})$ $\mathcal{O}(kr+dr)$
LoRA-Rite [75] $\mathcal{O}(kr^{2}+dr^{2})$ $\mathcal{O}(kr+dr)$
LoRA-Pro [62] $\mathcal{O}(kdr)$ $\mathcal{O}(kd)$
AltLoRA $\mathcal{O}(kr^{2}+dr^{2})$ $\mathcal{O}(kr+dr)$
Table 1: Comparison with Existing Work
Our main contributions are summarized as follows:
  • We propose AltLoRA, a novel PEFT method that efficiently approximates the full gradient via alternating projections onto the low-rank subspaces $A$ and $B$. Moreover, we design a new momentum mechanism that operates within LoRA’s memory constraints, enabling effective optimization of momentum within the low-rank space.
  • Theoretically, we prove that AltLoRA ensures stable feature learning in the infinite-width neural network regime and, more generally, maintains transformation invariance, even when incorporating momentum. We also provide convergence guarantees for fine-tuning overparameterized two-layer ReLU networks.
  • Empirically, we show the effectiveness of AltLoRA through extensive experiments on tasks including natural language understanding, dialogue generation, mathematical reasoning, and code generation. AltLoRA consistently outperforms existing LoRA-based methods.
2 Related Work
Low-rank adaptation(LoRA)([23]) has been the subject of extensive research in foundation models([48, 4, 1, 31, 51, 57]), with numerous variations and improvements ([32, 26, 30, 74, 11, 25, 87]). One line of research focuses on dynamically adjusting the LoRA rank during training. This includes DyLoRA[58], IncreLoRA[78], and AdaLoRA[82]. Another line of work involves enhancing LoRA performance through the addition of extra scaling matrices, which include DoRA[39] and DeepLoRA[74]. These directions are orthogonal to our work. Regarding the optimization of LoRA, we find that the following topics are close to our work.
Stable Feature Learning Under the infinite-width NN setting([17, 52]), LoRA+([19]) finds that the standard LoRA is inefficient and they propose to use different learning rates for $A$ and $B$. To provide a careful choice of hyperparameters for efficient use of LoRA, a line of work analyzes LoRA under efficient learning ([18, 77]). Noticeably, [77] introduces preconditioners under a Riemannian metric ([44]) and updates LoRA by using scaled gradients of $A$ and $B$ simultaneously. While their method aims to improve stability and efficiency, it is important to note that their goal is not to approximate the full gradient. This approach does not yield an optimal approximation to the full gradient update. Moreover, [75] proposes an adaptive matrix preconditioning method preserving transformation-invariant, a sufficient condition for stable feature learning.
Approximation full-tuning or full gradient To fill the gap between LoRA and full fine-tuning, there are two lines of work with different motivations. The first class of work focuses on the initialization, like [62]. It proposes to make the initialization of LoRA align with the full-finetuning directly. However, after the first step, how difference between LoRA and full-tuning is unknown. The second line of work focuses on optimizing LoRA properly over the optimization trajectory([62, 47, 83]). Noticeably, [62] proposes to optimize the gradients of $A$ and $B$ together to approximate the full gradient. But the optimal approximation is hard to find under practical conditions and aligning momentum towards the full gradient requires storing a full-size matrix ($k\times d$) in their algorithm. These challenges also exist in later work ([47]).
Gradient Projection in LoRA Motivated by the view that LoRA updates can be viewed as performing random projection from the full gradient, F-LoRA([16]) achieves high-rank updates by resampling the projection matrices. There are also some approaches that propose training networks with low-rank factorized weights from scratch ([60, 30]). Random projection is also applied in Ga-LoRA([84]) and following work([37, 8]), but they need to access the full model and can’t store the low-rank adapter in the end. On the contrary, without full-parameter learning, we use gradient projection to keep the gradient best preserved in the low-rank spaces.
Alternating Update To the best of our knowledge, we haven’t found the existing work of updating LoRA alternately in the centralized setting, but in the decentralized setting, i.e., Federated Learning, we notice [6] used the alternating strategy to address the challenge of inaccurate model aggregation([64, 3, 54]) with computational and communication efficiency. Besides, in the centralized setting, [81] proposes to freeze $A$ and update $B$, which would be regarded as a specific case of our work to do alternating minimization.
Scaled Gradient Descent Our proposed methods are also closely related to scaled gradient descent(Scaled GD) in traditional low-rank matrix estimation under over-parameterization and ill-conditioning ([55, 56, 27, 43]). Notably, [55] shows that the scaled GD would keep the convergence independent of the condition number. Different variants of scaled GD have been proposed and studied in work ([69, 79, 9, 80]). For the alternating scaled GD, [28] finds that it would enable faster convergence with larger step sizes compared with scaled GD. And [40] provably shows that alternating scaled GD would achieve a linear convergence rate, starting from arbitrary random initialization.
3 Preliminary
Let us first revisit the optimization paradigm of LoRA [23]. If we denote the loss function as $L$, i.e., $L(A,B):=L(W+sBA)$, we can derive the gradient w.r.t $A$ and $B$ as follows:
$\nabla_{A}L:=\frac{\partial L}{\partial A}=\frac{\partial L}{\partial W}\frac{\partial W}{\partial A}=sB^{T}(\nabla_{W}L),\quad\nabla_{B}L:=\frac{\partial L}{\partial B}=\frac{\partial L}{\partial W}\frac{\partial W}{\partial B}=s(\nabla_{W}L)A^{T}.$ (1)
Here, as the full gradient is multiplied by the low-rank matrices to constitute the gradient of LoRA, it implicitly compresses the full gradient into the low-rank spaces. Suppose we use gradient descent to update $A$ and $B$, then the model parameter in the $(t+1)$-th iteration is:
$\begin{split}W_{t+1}&=W_{0}+sB_{t+1}A_{t+1}\\ &\approx W_{0}+sB_{t}A_{t}-s\eta(\nabla_{B_{t}}L)A_{t}-s\eta B_{t}(\nabla_{A_{t}}L)\\ &=W_{t}-s\eta(\nabla_{B_{t}}L)A_{t}-s\eta B_{t}(\nabla_{A_{t}}L).\end{split}$ (2)
Here, we omit the term related to $\eta^{2}.$ Compared with the full gradient update $-\eta\nabla_{W}L$, LoRA’s gradient can approximate the full gradient as long as $sB(\nabla_{A}L)+s(\nabla_{B}L)A$ is close to $\nabla_{W}L$. With a similar motivation, some previous work analyzes the approximation based on the Frobenius norm ([61, 62, 47]). Noticeably, LoRA-Pro [62] achieves gradient approxiation by adjusting the gradients of matrices $A$ and $B$ based on the following solutions:
$g^{A}=\frac{1}{s}(B^{T}B)^{-1}B^{T}(\nabla_{W}L)+XA,g^{B}=\frac{1}{s}[I-B(B^{T}B)^{-1}B^{T}](\nabla_{W}L)A^{T}(AA^{T})^{-1}-BX,$ (3)
where $X\in\mathbb{R}^{r\times r}$ denotes an ancillary matrix and its selection is crucial and challenging for LoRA-Pro. As shown in their ablation studies, the selection of $X$ would vary the performance of the evaluation significantly. Besides, to obtain a unique solution for $X$, LoRA-Pro imposes additional uncommon assumptions to solve a Sylvester equation. However, even selecting a unique $X$, the equivalent gradient($sBg^{A}+sg^{B}A$) established by LoRA-Pro is independent of $X$, which implies that $X$ is only used to distinct the gradient of $A$ and $B$ when jointly updating and doesn’t influence the model update. It motivates the development of a more efficient alternating and eliminates the influence of $X$. To circumvent the ambiguity and inefficiency introduced by this joint updating strategy, we propose an alternating update strategy that approximates the full gradient as long as $sB(\nabla_{A}L)$ or $s(\nabla_{B}L)A$ is close to $\nabla_{W}L$.
[…]
4 Methodology
4.1 Alternately Approximating the Full Gradient via Low-Rank Adaptation
We propose an alternating update scheme, where we update $A$ first and then update $B$ based on the new $A$. Define the low-rank modules as $A_{t}$ and $B_{t}$ at the $t$-th iteration, and the approximated gradients as ~$\tilde{\nabla}_{A}L$ and ~$\tilde{\nabla}_{B}L$, respectively. We begin by obtaining the optimal scaling gradient of $A$ by solving
$\min_{\tilde{\nabla}_{A_{t}}L}\|sB_{t}(\tilde{\nabla}_{A_{t}}L)-\nabla_{W_{t}}L\|^{2}_{F},$ (4)
where $\|\cdot\|^{2}_{F}$ denotes the Frobenius norm squared—sum of squares of all entries in the matrix. Then by gradient descent, we can update $A$ and the full model as
$\begin{split}A_{t+1}\leftarrow A_{t}-\eta\tilde{\nabla}_{A_{t}}L,\quad W_{t+\frac{1}{2}}\leftarrow W_{t}-\eta B_{t}(\tilde{\nabla}_{A_{t}}L),\end{split}$ (5)
where we update the full model at $(t+1/2)$-th iteration to keep consistent with the joint update [62] (update $A$ and $B$ in one iteration). In our experiment, without any ambiguity, we treat the update $A$ or $B$ as a single step (see Algorithm 1). After doing backpropagation w.r.t $A$, the gradient of $B$ doesn’t approximate the full gradient at time $t$ since the full model has been update to the state of $(t+1/2)$. Then we minimize the discrepancy between the full gradient at $W_{t+\frac{1}{2}}$ and the approximating gradient constructed by $B_{t}$ as follow
$\min_{\tilde{\nabla}_{B_{t}}L}\|s(\tilde{\nabla}_{B_{t}}L)A_{t+1}-\nabla_{W_{t+\frac{1}{2}}}L\|_{F}^{2}.$ (6)
Then by gradient descent, we can update $B$ and the full model as
$\begin{split}B_{t+1}\leftarrow B_{t}-\eta\tilde{\nabla}_{B_{t}}L,\quad W_{t+1}\leftarrow W_{t+\frac{1}{2}}-\eta(\tilde{\nabla}_{B_{t}}L)A_{t+1}.\end{split}$ (7)
The following theorem gives the closed-form solution of Problems (4) and (6).
Theorem 1.
Assume $B_{t}\in\mathbb{R}^{k\times r}$ and $A_{t}\in\mathbb{R}^{r\times d}$ are full rank for any $t$, i.e. rank($B_{t}$) = rank($A_{t}$) = $r$. Solving Problems (4) and (6) yields the unique closed-form solutions
~$\begin{split}\tilde{\nabla}_{A_{t}}L&=\frac{1}{s}(B_{t}^{T}B_{t})^{-1}B_{t}^{T}(\nabla_{W_{t}}L)=\frac{1}{s^{2}}(B_{t}^{T}B_{t})^{-1}\nabla_{A_{t}}L\\ \tilde{\nabla}_{B_{t}}L&=\frac{1}{s}\left(\nabla_{W_{t+\frac{1}{2}}}L\right)A_{t+1}^{T}(A_{t+1}A_{t+1}^{T})^{-1}=\frac{1}{s^{2}}\nabla_{B_{t}}L(A_{t+1}A_{t+1}^{T})^{-1},\end{split}$ (8)
where $\nabla_{A_{t}}L$ and $\nabla_{B_{t}}L$ are the gradients of LoRA defined in Equation (1).
Theorem 1 shows that both problems admit unique optimal solutions for ~$\tilde{\nabla}_{A_{t}}L$ and ~$\tilde{\nabla}_{B_{t}}L$, which only requires full rank. Therefore, it offers a new gradient approximation with less computational cost and promotes a more efficient updating strategy. Besides, instead of accessing the full gradient like full fine-tuning, the optimal gradient approximation only requires the standard gradient of $A$ or $B$ by backpropagation at each step and calculating the inverse of a small matrix with size $r\times r$.
Theorem 1 requires that the matrix $B_{t}$ and $A_{t}$ are full rank, but in the over-parameterized cases, the assumption is hard to achieve. To alleviate it, if we penalize the Frobenius norm of these two approximated gradients, i.e., weight decay, the condition can be eliminated (see Corollary 1 in Appendix). For simplicity, in the rest of the paper, we focus on the modified gradient in (8) for analysis. The closed-form solution in (8) yields the following full model update(with gradient descent)
$\begin{split}W_{t+1}&=W_{t+\frac{1}{2}}-\eta(\tilde{\nabla}_{B_{t}}L)A_{t+1}\\ &=W_{t+\frac{1}{2}}-\eta(\nabla_{W_{t+\frac{1}{2}}}L)A_{t+1}^{T}(A_{t+1}A_{t+1}^{T})^{-1}A_{t+1}\\ &=W_{t}-\eta B_{t}\tilde{\nabla}_{A_{t}}L-\eta(\nabla_{W_{t+\frac{1}{2}}}L)A_{t+1}^{T}(A_{t+1}A_{t+1}^{T})^{-1}A_{t+1}\\ &=W_{t}-\eta B_{t}(B_{t}^{T}B_{t})^{-1}B_{t}^{T}(\nabla_{W_{t}}L)-\eta(\nabla_{W_{t+\frac{1}{2}}}L)A_{t+1}^{T}(A_{t+1}A_{t+1}^{T})^{-1}A_{t+1}\\ &=W_{t}-\eta Proj_{c(B_{t})}(\nabla_{W_{t}}L)-\eta(\nabla_{W_{t+\frac{1}{2}}}L)Proj_{r(A_{t+1})}.\end{split}$ (9)
Interestingly, the proposed solution for gradient approximation in (8), is consistent with the literature work [55, 79, 69, 28, 40] called scaled gradient descent [43, 44] in low-rank matrix estimation [50]. Therefore, the view of gradient approximation would provide a novel interpretation of applying scaled gradient descent within the broader context of low-rank matrix decomposition. As optimizing LoRA with momentum for acceleration is a standard way in the literature [7, 23, 19], we will discuss how to properly design momentum within the low-rank space inspired by gradient approximation.
4.2 Proper Momentum Design within the Low-Rank Subspaces
For LoRA [23] and its variants [19, 82] without allowing full-parameter learning, the parameterization restricts both the gradient and the momentum updates to low-rank subspaces as the memory cost is $\mathcal{O}(kr+dr)$. As we have shown, the optimal gradient approximation under this constraint is obtained by projecting the full gradient onto the low-rank subspace. This insight naturally motivates the need to also align the momentum optimally within the same low-rank space, in order to fully leverage momentum-based acceleration under low-rank constraints.
Since the momentum evolves throughout training, it is essential to dynamically optimize it. For simplicity, we focus on the optimization paradigm for $B$ and develop our method inductively. Given the aligned momentum $M_{t}^{B}$ within the low-rank space $A_{t}$ at time $t$, the alternating update strategy proceeds by updating $A$ to $A_{t+1}$ and then aligning $M_{t}^{B}$ with the new low-rank space $A_{t+1}$. To this end, we first recover $M_{t}^{B}$ to the full-dimensional space, and then project it onto the new subspace spanned by $A_{t+1}$, like gradient approximation. The following theorem formalizes this key idea.
Theorem 2.
Assume $A_{t+1}A_{t+1}^{T}$ is full-rank, i.e., rank$(A_{t+1}A_{t+1}^{T})$ = $r$. If $M_{t}^{B}$ has aligned with the low-rank space $A_{t}$ in the $t$-th iteration, by minimizing the following problem
$\begin{split}\min_{\tilde{M}_{t}^{B}}\|M_{t}^{B}A_{t}-\tilde{M}_{t}^{B}A_{t+1}\|^{2}_{F}.\end{split}$ (10)
We can find $\tilde{M}_{t}^{B}=M_{t}^{B}A_{t}A_{t+1}^{T}(A_{t+1}A_{t+1}^{T})^{-1}$, which makes the momentum aligned with the new low-rank space $A_{t+1}$ optimally.
Theorem 2 shows that it’s only necessary to store two small matrices so that we can optimize momentum properly. Similar to Section 4.1, we can also remove the assumption of full rank here (see Corollary 2 in Appendix). In contrast to LoRA-Pro with full-parameter learning (Space Complexity $\mathcal{O}(kd)$), we aim to strictly satisfy the space complexity $\mathcal{O}(kr+dr)$ for parameter efficiency and keep mentum adaptively aligned with the low-rank spaces as gradient approximation does.
A similar notion of momentum design is explored in [16, 20], where down-projection and up-projection matrices are employed to transfer compressed gradients across low-rank spaces. In contrast, we derive the optimal alignment directly within the low-rank subspaces to preserve gradient information. In Section 5.2, we theoretically demonstrate that aligning momentum with the low-rank space guarantees transformation invariance, whereas LoRA [23] and its variants [19, 82] have misaligned momentum undermining this robustness [75].
After analyzing how to efficiently optimize both the gradient and momentum under limited resource constraints, we summarize our proposed algorithm, AltLoRA, in Algorithm 1. Unlike the joint update strategy, AltLoRA updates only one of the low-rank matrices, either $A$ or $B$, at each step, based on the scaled gradient and momentum presented in Theorems 1 and 2. The number of trainable parameters at each step is reduced by half compared to the joint update. Designed as a practical PEFT method, AltLoRA can be seamlessly integrated into existing libraries such as Hugging Face [65] (see Appendix C.1 for implementation details). To further accelerate and stabilize the training paradigm of AltLoRA, we introduce AltLoRA+, an enhanced variant that naturally incorporates second-moment estimates similar to AdamW (see Algorithm 2 for details).
Input: Momentum states $M_{0}^{A}$, $M_{0}^{B}$; scaling factor $s=\frac{\alpha}{r}$; learning rate $\eta$; momentum coefficient $\beta_{1}$; total steps $T$; weight decay $\gamma$
Output: Final matrices $A_{T}$ and $B_{T}$
1 for $t=0,\dots,T-1$ do
2       if $t\bmod 2=0$  then
3             Update $A$:
4             Only backpropagate w.r.t. $A_{t}$ and obtain $\nabla_{A_{t}}L$
5             ~$\tilde{\nabla}_{A_{t}}L=\frac{1}{s^{2}}(B_{t}^{\top}B_{t})^{-1}\nabla_{A_{t}}L$
6             $\tilde{M}_{t}^{A}=(B_{t}^{\top}B_{t})^{-1}B_{t}^{T}B_{t-1}M_{t-1}^{A}$
7             $M_{t}^{A}\leftarrow\beta_{1}\tilde{M}_{t}^{A}+(1-\beta_{1})\tilde{\nabla}_{A_{t}}L$
8             $A_{t+1}\leftarrow A_{t}-\eta(M_{t}^{A}+\gamma A_{t})$
9
10      else
11             Update $B$:
12             Only backpropagate w.r.t. $B_{t}$ and obtain $\nabla_{B_{t}}L$
13             ~$\tilde{\nabla}_{B_{t}}L=\frac{1}{s^{2}}\nabla_{B_{t}}L(A_{t+1}A_{t+1}^{\top})^{-1}$
14             $\tilde{M}_{t}^{B}=M_{t-1}^{B}A_{t}A_{t+1}^{\top}(A_{t+1}A_{t+1}^{\top})^{-1}$
15             $M_{t}^{B}\leftarrow\beta_{1}\tilde{M}_{t}^{B}+(1-\beta_{1})\tilde{\nabla}_{B_{t}}L$
16             $B_{t+1}\leftarrow B_{t}-\eta(M_{t}^{B}+\gamma B_{t})$
17
18
Algorithm 1 AltLoRA: Gradient Approximation via Alternating Projection with Proper Momentum Design under LoRA’s Memory Constraint
Time Complexity and Space Complexity. When $r\ll min\{k,d\}$, the time and memory cost of AltLoRA and AltLoRA+ is similar to the standard LoRA and more efficient compared with LoRA-Pro. The additional computational cost takes $\mathcal{O}(r^{3})$ time, and since $r$ is very small, this overhead is negligible when compared with the back-propagating time. In the experiment, we will show that the delay time compared with LoRA is mild even when the rank $r$ increases. (see Table 3).
5 Theoretical Analysis
5.1 Stable Feature Learning
Given the current trend of increasing model sizes ([72, 45, 71]), it raises a lot of attention to analyze the asymptotic training behavior of neural networks as the number of neurons approaches infinity ([52, 17, 70]). There is a line of work in LoRA ([19, 18, 77]) considering the infinite-width NN setting. To achieve stable feature learning (see Definition 1 in Appendix D.1), they propose a fine-grained choice of hyperparameters in the original LoRA, like the learning rate [19], the initialization ([18]), and the optimizer ([77]). The core idea is that the update increment over the loss function or parameter should be of constant magnitude, which ensures that neither the NN predictions nor the increments explode or vanish as the NN size increases, thereby leading to stable training dynamics. First, we demonstrate that our method achieves stable feature learning on a toy model in Appendix D.1.1. We then prove that this stability extends to arbitrary LoRA ranks and holds for AltLoRA and AltLoRA+, which we formalize in the theorem below. For clarity of presentation, we omit the scaling factor $s$ in the subsequent theorems and analysis.
Theorem 3 (Informal).
Assume that, with the input $x$, $BAx$ has dimension $\mathcal{O}(n)$. In Algorithm 1 or Algorithm 2, if we use the same learning rate $\eta=\mathcal{O}(1)$ to update $A$ and $B$, it would achieves stable feature learning. Moreover, without momentum in AltLoRA or AltLoRA+, the model update achieves stable feature learning as well with
$W_{t+1}=W_{t}-\eta Proj_{c(B_{t})}(\nabla_{W_{t}}L)-\eta(\nabla_{W_{t+\frac{1}{2}}}L)Proj_{r(A_{t+1})},$ (11)
where
$\eta Proj_{c(B_{t})}(\nabla_{W_{t}}L),\eta(\nabla_{W_{t+\frac{1}{2}}}L)Proj_{r(A_{t+1})}\in\mathcal{O}(1).$
However, when doing joint update ([77]), the update will introduce additional across term
$\eta^{2}(\nabla_{W_{t}}L)A_{t}^{T}(A_{t}A_{t}^{T})^{-1}(B_{t}^{T}B_{t})^{-1}B_{t}^{T}(\nabla_{W_{t}}L)\in\mathcal{O}(1).$
The across term is indeed the second order term w.r.t $\eta$, but it is same magnitude as $\eta(\nabla_{W_{t}}L)Proj_{r(A_{t})}$ and $\eta Proj_{c(B_{t})}(\nabla_{W_{t}}L)$ in infinite-width NN setting.
In Theorem 3, AltLoRA and AltLoRA+ achieve stable feature learning. Moreover, as the joint update would introduce the cross term with an unignorable magnitude (especially $\eta$ is $\mathcal{O}(1)$ instead of $\mathcal{O}(1/n$) in the toy model), joint update with scaled gradient descent ([77]) breaks the clean interpretation of projecting the full gradient onto low-rank subspaces and degrade the performance as our experiment studies show later.
5.2 Transformation Invariance
With the motivation that an optimizer should yield the same update to the full model regardless of the specific factorization, transformation invariance, as a sufficient condition for stable feature learning, is proposed by LoRA-RITE [75]. Here, we will prove that our designed gradient and momentum in Algorithm 1 would be inherently robust as transformation invariance.
Definition 1.
If there are two pairs of LoRA matrix $(A_{1},B_{1}),(A_{2},B_{2})$ can represent the same finetuned weight $W=W_{0}+B_{1}A_{1}=W_{0}+B_{2}A_{2}$. An optimizer exhibits transformation invariance if its updates, $(\delta A_{1},\delta B_{1})$ and $(\delta A_{2},\delta B_{2})$ satisfy
$\begin{split}&W_{0}+(B_{1}+\delta B_{1})(A_{1}+\delta A_{1})=W_{0}+(B_{2}+\delta B_{2})(A_{2}+\delta A_{2})\\ &\Rightarrow(B_{1}+\delta B_{1})(A_{1}+\delta A_{1})=(B_{2}+\delta B_{2})(A_{2}+\delta A_{2}).\end{split}$ (12)
LoRA-RITE [75] notices that, after combining scaled gradient descent with element-wise Adam in [77], the ScaledAdam can’t preserve transformation invariance. As the momentum is optimized properly, we will analyze how AltLoRA keeps transformation invariance naturally, especially when incorporating momentum.
Recall the definition of projection matrices in Equation (9): $Proj_{c(B_{t})}:=B_{t}(B_{t}^{T}B_{t})^{-1}B_{t}^{T}$ (or $Proj_{r(A_{t})}:=A_{t}^{T}(A_{t}A_{t}^{T})^{-1}A_{t}$). The following lemma provides insight into how Algorithm 1 achieves transformation invariance.
Lemma 1.
If any two pairs of LoRA factors$(A_{1},B_{1}),(A_{2},B_{2})$ satisfying
$W=W_{0}+B_{1}A_{1}=W_{0}+B_{2}A_{2},$ (13)
then $\quad\quad\quad\quad\quad\quad Proj_{c(B_{1})}=Proj_{c(B_{2})},\quad Proj_{r(A_{1})}=Proj_{r(A_{2})}$
Even though the full model update can be decomposed into different pairs of low-rank adaptations, within each pair of LoRA factors, the column space of $B$ (or the row space of $A$) is equivalent to the column space (or the row space) of the full model update. Therefore, the projection matrix would be preserved invariant over the pairs of low-rank adaptation.
Theorem 4.
AltLoRA in Algorithm 1 is transformation-invariant.
Building on the insight from Lemma 1, we leverage the invariance of the projection matrix to the low-rank subspaces to approximate the full gradient via the gradient and moment information. As a result, with the goal of gradient approximation without full-parameter learning, our method achieves transformation invariance inherently. LoRA-RITE [75] is also aware of the equivalence of low-rank spaces, but they do not notice or exploit the invariance of the projection matrix. Instead, they design an unmagnified gradient requiring polar decomposition at each iteration, which introduces additional computational overhead. In contrast, our method avoids polar decomposition, contributing to its superior efficiency (see Table 3). LoRA-Pro [62] also achieves transformation invariance but does so without adhering to LoRA’s memory constraint. AltLoRA in Algorithm 1, by comparison, strictly follows the memory budget of LoRA while preserving transformation invariance through a more efficient design. While Algorithm 2 does not currently maintain transformation invariance under second-order momentum, this opens an exciting avenue for future research. In Appendix D.2, we provide a detailed discussion on why extending our first-order momentum design to the second order poses fundamental challenges. Despite this, AltLoRA+ achieves substantial empirical gains over LoRA and its variants, demonstrating the practical strength of our approach even when we only keep the transformation-invariant up to the second momentum.
5.3 Convergence Analysis
Following [77], we provide a convergence analysis of AltLoRA (or AltLoRA+) without momentum within the over-parameterized two-layer ReLU NN tuning problem (see Appendix D.3). We show that the convergence is independent of the condition number of the data matrix. In contrast to [77], we impose fewer assumptions to establish the convergence analysis. Notably, we don’t require the extended spectral initialization in Definition 7.3 [77]. In our experimental study, AltLoRA (AltLoRA+) can achieve superior performance with the variant of initialization used by LoRA and its variants (see Appendix E.3.2), which supports our insight empirically.
6 Experimental Results
This section empirically shows the effectiveness of our approach across various model architectures and datasets. Section 6.1 summarizes the experimental settings and results on supervised fine-tuning (SFT) benchmark tasks, and Section 6.2 provides details of the setup and results for natural language understanding tasks. Finally, ablation studies from multiple perspectives are presented in Section 6.3. The code for our project is available at https://anonymous.4open.science/r/AltLoRA-DB7C.
6.1 Experiments on SFT of LLM: Natural Language Generation
Training Details. We assess our methods on dialogue generation with the WizardLM dataset [68], mathematical reasoning with the MetaMathQA dataset [76], and code generation with the CodeFeedBack dataset [86] using the LLama-3.1-8B and Llama-3-8B models [15] (see Appendix E.1). We compare AltLoRA and AltLoRA+ with the pretrained model, full fine-tuning, LoRA [23], PisSSA[42], rsLoRA[29], LoRA+[19], DoRA[39], AdaLoRA[82], LoRA-GA[61], LoRA-Rite [75]and LoRA-Pro[62]. To ensure fair comparisons, we closely follow the experimental protocol established by [62]. Unless otherwise stated, we fine-tune models using default hyperparameters (if used): $\beta_{1}=0.9$, $\beta_{2}=0.999$, and zero weight decay. We adopt a cosine learning rate schedule with a warm-up ratio of 0.03. LoRA adapters are applied to $\{Q,K,V,O\}$ layers. By default, we set the rank to $r=8$ and the scaling factor to $\alpha=32$ for dialogue generation tasks, and $r=8$, $\alpha=16$ for the mathematical reasoning and code generation tasks. We carefully grid search the learning rates  [^1] See Appendix E.1 for details of learning rate grid search. We set the sequence length to 1024 and the macro batch size to 4 for math and code tasks, and macro batch size to 8 for dialogue generation. All experiments are conducted on NVIDIA A100 and NVIDIA A6000 GPUs.. To obtain a reliable estimate of model performance, we perform three runs with different random seeds and report the average and standard deviation of the results.
Evaluations. We evaluate the baselines similar to [62]. Specifically, for the dialogue generation task, we use the MT-Bench dataset [85] with GPT-4o, with scores ranging from 1 to 10. We report the score from the first turn as our metric. For the math task, we evaluate the model on the GSM8K test set [10] using the LLM Evaluation Harness [14], and we report the exact match accuracy. For the code generation task, we evaluate on the HumanEval dataset [5] and report the PASS@1 metric.
Results. Table 2 presents our experimental results, which demonstrates AltLoRA superior performance. With a rank of 8, AltLoRA achieves noticeable improvement over the original LoRA: 0.5 on MT-bench, 8.38 on GSM8K and 3.1 on HumanEval using Llama-3.1-8B. Notably, AltLoRA achieves significantly higher scores on MT-Bench compared to LoRA-Pro and Full FT. In addition, AltLoRA+ yields improvements over LoRA-Pro on both GSM8K and HumanEval, and AltLoRA+ obtains better performance in mathematical reasoning than Full FT. These further demonstrate the effectiveness of the new design gradient and momentum. The additional study on Llama-3-8B model (see Table 8 in Appendix E.1) also demonstrates a clear advantage over baseline methods.
Method MT-Bench GSM8K HumanEval
PreTrain 5.93$\pm$0.08 51.34$\pm$1.38 36.15$\pm$1.97
Full FT 6.31$\pm$0.04 73.31$\pm$0.32 50.81$\pm$1.10
LoRA 6.06$\pm$0.02 66.11$\pm$1.43 40.31$\pm$1.34
PiSSA 5.15$\pm$0.10 67.78$\pm$1.11 42.44$\pm$1.11
rsLoRA 6.10$\pm$0.06 68.12$\pm$0.44 43.91$\pm$1.44
LoRA+ 6.40$\pm$0.06 72.33$\pm$1.33 44.10$\pm$1.38
DoRA 6.08$\pm$0.03 68.33$\pm$0.88 42.13$\pm$1.31
AdaLoRA 6.08$\pm$0.05 72.63$\pm$1.45 42.21$\pm$2.66
LoRA-GA 6.00$\pm$0.09 70.33$\pm$0.91 42.01$\pm$1.21
LoRA-Pro 6.19$\pm$0.03 73.12$\pm$0.56 43.13$\pm$1.45
LoRA-Rite 6.10$\pm$0.01 74.10$\pm$0.31 43.12$\pm$0.51
AltLoRA 6.56$\pm$0.04 74.49$\pm$0.57 45.91$\pm$1.14
AltLoRA (rank=32) 6.39$\pm$0.04 73.24$\pm$0.29 46.87$\pm$1.49
AltLoRA (rank=128) 6.27$\pm$0.01 74.11$\pm$0.21 45.41$\pm$1.65
AltLoRA+ 6.16$\pm$0.02 76.91$\pm$0.31 50.10$\pm$1.35
AltLoRA+ (rank=32) 6.10 $\pm$0.02 76.32$\pm$0.29 49.97$\pm$1.52
AltLoRA+ (rank=128) 6.07$\pm$0.03 77.08$\pm$0.83 49.77$\pm$1.58
Table 2: Comparison of different LoRA variants on MT-Bench, GSM8K, and HumanEval benchmarks on Llama-3.1-8B-Base. Bold indicates the best result, underline represents the second-best one.
Table 3: Comparison of memory usage and training time across different fine-tuning methods.
Method Memory Cost Training Time
Full FT $>48$ GB 4h 23min
LoRA 22.26 GB 2h 13min
LoRA-Rite 25.39 GB 2h 44min
LoRA-Pro 40.12 GB 4h 5min
AltLoRA 22.56 GB 2h 34min
AltLoRA(rank=32) 23.11 GB 2h 41min
AltLoRA(rank=128) 25.11 GB 2h 52min
AltLoRA+ 23.16 GB 2h 38min
AltLoRA+(rank=32) 24.98 GB 2h 45min
AltLoRA+(rank=128) 27.76 GB 2h 56min
Memory and Time Consumptions. In Table 3, we also compare the memory cost and training time of our methods with Full FT, LoRA, LoRA-Rite and LoRA-Pro on Llama-3.1-8b mode. Without full-parameter learning, we have a comparable memory cost and training time close to LoRA. After taking a higher rank of LoRA, the memory cost and computation cost won’t increase significantly. However, as LoRA-Pro requires storing the full size first-order momentum and second-order momentum, it leads to an unignorable cost like Full FT. As LoRA-Rite incurs additional calculations like polar decomposition, it also increase the computation time.
6.2 Experiments on Natural Language Understanding
Training and Evaluation Details. We assess our methods natural language understanding on a subset of GLUE benchmark dataset with fine-tuning a T5-base[49] model. We compare AltLoRA and AltLoRA+ with the full fine-tuning, LoRA [23], PisSSA[42], rsLoRA[29], LoRA+[19], DoRA[39], AdaLoRA[82], LoRA-GA[61], and LoRA-Pro[62]. We fine-tune the T5-based model [49] with our methods and the baselines on a subset of GLUE datasets [59]: MNLI, SST2, CoLA, QNLI, and MRPC. We use the accuracy as the evaluation metric. To ensure fair comparison, all experiments are run three times with different random seeds, and we report the mean and standard deviation of the results. Due to space constraints, additional experimental details are provided in Appendix E.1.
Results. As shown in Table 4, AltLoRA+ outperforms the baselines on average. In particular, it achieves the highest score on MRPC, the second-highest on CoLA, MNLI, and SST-2 datasets.
Table 4: Performance of fine-tuning T5-Base on 5 sub-tasks of the GLUE benchmark. Bold indicates the best result, underline represents the second-best one, and * marks results reported from [61].
Method MNLI SST-2 CoLA QNLI MRPC Average
Full 86.29$\pm$0.01 93.97$\pm$0.06 80.87$\pm$0.05 93.02$\pm$0.03 86.89$\pm$0.13 88.21
LoRA 85.32$\pm$0.01 93.76$\pm$0.05 81.31$\pm$0.20 92.96$\pm$0.09 86.03$\pm$0.24 87.88
RSLoRA 85.23$\pm$0.01 93.96$\pm$0.06 81.21$\pm$0.14 93.12$\pm$0.09 86.27$\pm$0.24 87.96
DoRA 85.58$\pm$0.03 93.65$\pm$0.06 81.16$\pm$0.04 93.04$\pm$0.06 86.14$\pm$0.12 87.91
LoRA+ 85.32$\pm$0.06 93.92$\pm$0.11 81.21$\pm$0.06 92.97$\pm$0.03 86.25$\pm$0.16 87.93
PiSSA 85.87$\pm$0.04 93.84$\pm$0.06 81.90$\pm$0.05 93.16$\pm$0.09 86.64$\pm$0.12 88.28
LoRA-GA 85.70$\pm$0.09 94.11$\pm$0.18 80.57$\pm$0.20 93.18$\pm$0.06 85.29$\pm$0.24 87.77
AdaLoRA 85.45$\pm$0.11 93.92$\pm$0.09 80.31$\pm$0.05 91.66$\pm$0.05 86.16$\pm$0.60 87.50
LoRA-Pro 85.70$\pm$0.11 93.92$\pm$0.10 78.42$\pm$0.03 93.15$\pm$0.03 86.54$\pm$0.50 87.55
AltLoRA 85.26$\pm$0.04 93.87$\pm$0.05 80.44$\pm$0.09 91.56$\pm$0.01 86.60$\pm$0.99 87.55
AltLoRA+ 85.81$\pm$0.03 94.03$\pm$0.12 81.44$\pm$0.30 92.99$\pm$0.03 87.25$\pm$1.12 88.30
6.3 Ablation Study
Figure 1 presents an ablation study of the learning rate $\eta$ and the scaling factor $\alpha$ for LoRA, AltLoRA and AltLoRA+, using the LLaMA 3.1-8B model on mathematical reasoning tasks. The results show that our proposed methods are robust in learning rate and the scaling factor with consistent superior performance. Moreover, it shows that $\alpha=16$ obtains overall better performance compared to $\alpha=8$ and $\alpha=32$. The influence of increasing rank is reported in Table 2.
Figure 1: Evaluation Accuracy of LoRA, AltLoRA and AltLoRA+ for various learning rate $\eta$ and scaling factor $\alpha$ combination on the GSM8K datasets using Llama-3.1-8B.
Next, to evaluate the effectiveness of alternating strategies, we compare them against the joint update method.
Ablation study on the updating strategy In Table 5, in contrast to joint update with scaled gradient descent [77], AltLoRA can optimally approximates the full gradient with alternating update and obtain better performance in evaluation. Interestingly, we find that the alternating update scheme—where matrix $B$ is updated before $A$—consistently yields better performance. One possible explanation is that, under the standard initialization where $B$ is set to zero, updating $A$ first does not lead to meaningful descent.
GSM8K LoRA AltLoRA AltLoRA+
Alternating ($A$ first) 66.11 74.49 76.91
Alternating ($B$ first) 67.66 76.31 76.97
Joint Update 66.43 74.21 76.56
Table 5: Performance comparison of LoRA, AltLoRA and AltLoRA+ on the GSM8K and Llama 3.1 8B with different updating strategies.
As the approaches of multiple LoRA modules, such as in the mixture of LoRA experts, has gained popularity [35, 66], we also assess the impact of varying the number of experts in LoRA layers.
Ablation study on the number of LoRAs As low-rank adaptation comes to be a popular parameter-efficient technique for fine-tuning, it’s well applied to more complicated scenarios ([41, 73, 53, 21, 66]). Notably, a very significant application is to improve the structure of the mixture of experts with parameter efficiency([66, 35]), handling multiple tasks simultaneously ([46, 24]) and addressing catastrophic forgetting ([13]). Following the work ([66]), we explore the performance as the number of LoRAs varies and utilize the gating balancing loss. Additionally, we compare AltLoRA and standard LoRA on the GSM8K dataset using the Llama 3.1-8B model(see Table 6). In our experiments, the number of LoRA experts is set to $\{1,4,8\}$, and the entropy regularization weight is 0.0001. We observe that increasing the number of LoRA experts enhances the capacity of the language model, leading to improved performance.
Expert Num LoRA AdaLoRA LoRA+ AltLoRA AltLoRA+
1 66.11 72.63 72.33 74.49 76.91
4 67.43 71.71 71.27 75.01 77.33
8 67.89 70.34 71.44 75.33 76.94
Table 6: Comparison of the mixture of experts model, with different expert numbers on GSM8K and Llama 3.1-8B-Base
To further validate the robustness of our method with respect to initialization, as discussed in Section 5.3, we study different initialization strategies.
Ablation Study on Initialization. To further validate the robustness of our method with respect to initialization, as discussed in Section 5.3, we conduct an ablation study using different initialization strategies. "Gaussian" refers to the standard random initialization used in the original LoRA framework [23]. "Kaiming" denotes the widely adopted Kaiming initialization, which is designed to maintain variance stability across layers. "Spectral" represents an initialization strategy based on spectral decomposition, where we perform singular value decomposition (SVD) on the pretrained weight matrix and construct the low-rank components using the top-$r$ singular vectors, like the initialization proposed in [84]. In Table 7, we can see that with different initialization strategies, our method would achieve a superior performance over the standard LoRA. Without spectral initialization, using Kaiming initialization for $A$ and setting $B$ to be zero would achieve the best performance. Besides, to ensure the initial update of $BA$ is zero, one of the matrices must be initialized to zero. Notably, setting $B=0$ while using a small initialization for $A$ yields better performance compared to the reverse setup. This finding is consistent with observations in existing literature [18].
Initialization Strategy LoRA AltLoRA AltLoRA+
A B
Gaussian zero 66.37 73.13 76.87
zero Gaussian 66.18 72.13 76.50
Kaiming zero 65.11 74.49 76.91
zero Kaiming 67.10 74.03 76.88
Spectral zero 67.63 74.67 76.60
zero Spectral 67.10 74.61 76.37
Table 7: Comparison of the initialization strategies on GSM8K and Llama 3.1-8B-Base
These ablation studies collectively demonstrate that our method is robust to hyperparameter variations and is applicable to more complex model architectures.
6.3.1 Additional Ablation Study for Llama-3-8B Model
Method MT-Bench GSM8K HumanEval
PreTrain 5.63 49.96$\pm$ 0.38 34.76$\pm$0.37
LoRA 6.20 62.11$\pm$0.13 37.71$\pm$0.12
AltLoRA 6.05 64.39$\pm$0.23 40.81$\pm$0.47
AltLoRA+ 6.34 67.38$\pm$0.13 43.81$\pm$0.31
Table 8: Comparison of different LoRA variants on MT-Bench, GSM8K, and HumanEval benchmarks (accuracy in %) on Llama-3-8B-Base.
We further conduct ablation studies on the LLaMA 3-8B model to evaluate the robustness of our method under varying hyperparameter settings. As shown in Figure 2, we compare the performance of LoRA, AltLoRA, and AltLoRA+ on the GSM8K dataset across different learning rates and scaling factors $\alpha\in\{8,16,32\}$. AltLoRA+ consistently outperforms the baselines across all configurations, demonstrating both higher accuracy and stronger robustness to hyperparameter variation. We also have that all methods have better performance using $\alpha=16$.
[…]
In Table 8, we compare our method with existing approaches across the three tasks described on Llama-3-8B model. Our method further bridges the performance gap between LoRA and full fine-tuning.
7 Conclusion
We propose AltLoRA, a memory-efficient fine-tuning method that alternates updates of low-rank matrices to dynamically project both the gradient and momentum within low-rank subspaces. By leveraging an efficient closed-form gradient approximation and a principled momentum design, AltLoRA operates entirely under low-rank constraints while ensuring stable feature learning and transformation invariance without requiring full-parameter learning. Extensive experiments across diverse tasks demonstrate the superior performance of AltLoRA and its enhanced variant, AltLoRA+, over LoRA and its variants, narrowing the gap to full fine-tuning while retaining memory efficiency.